# Learning Equilibria from Data: Provably Efficient Multi-Agent Imitation Learning

### v1

@@ 6 Conclusion and Future Directions
This paper provides the first sample complexity analysis of behavioural cloning in the multi-agent setting. The provided upper bound depends on the single policy deviation concentrability coefficient, which is shown to be unavoidable in general. Unfortunately, it is quite easy to come up with MGs where the concentrability coefficient is unbounded. In this situation, we resort to expert queries and we introduce novel algorithms dubbed MAIL-BRO and MURMAIL, which achieve an $\varepsilon$-approximate Nash equilibrium with a polynomial number of expert queries and computational cost polynomial in all problem parameters. Several directions remain open. We outline a few of them in Appendix C.
Acknowledgments
Luca Viano is funded (in part) through a PhD fellowship of the Swiss Data Science Center, a joint venture between EPFL and ETH Zurich.
@@ Appendix A Notation
$\epsilon$-NE $\epsilon$-approximate Nash Equilibrium
@@
$N$ Number of trajectories in dataset
@@
$Error(\hat{\nu})$ Error term for player 2’s estimated policy
$Error(\hat{\mu})$ Error term for player 1’s estimated policy
@@
$\epsilon_{opt}$ Optimality gap for RL inner loop
@@ Appendix B Related Work
Strikingly, the scenario is completely swapped in the multi-agent setting. Our negative result Theorem 3.2 shows that given knowledge of the transition no significant improvements over BC can be expected. On the other side, our positive result Theorem 4.2 shows that in the interactive setting a consistent improvement is expected over BC if $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is large.
@@
Appendix C Future directions
We outline here few interesting future directions and research questions left open by our work.
Extension to deep imitation learning.
The current analysis is limited to tabular Markov Games. However, the main conceptual ideas easily carry on to deep imitation learning experiments. The largest theory-practice gap would be in the inner loop where UCBVI would need to be replaced by a Deep RL algorithm such as DQN Mnih et al. [2015] or Soft Actor Critic Haarnoja et al. [2018], just to name a few.
The focus of this work was to show the first sample complexity bound for a computationally efficient algorithm in the queriable expert setting. For the sake of simplicity, we did not try to optimize the dependence of the upper bound in the accuracy parameters $\varepsilon$, effective horizon $(1-\gamma)^{-1}$, states and actions cardinality $\left|{{\mathcal{S}}}\right|$ and $\left|{{\mathcal{A}}}\right|$. A possible direction of improvement is to derive a tighter analysis of the outer loop using faster rates for the regret of the squared loss ( see for example [Cesa-Bianchi and Lugosi, 2006, Chapter 3]).
Moreover, the upper bound could be improved by removing the need of the RL inner loop in MURMAIL. In general, replacing it with a no regret learner that minimizes the regret in an MDP with changing reward function and transitions is not possible because of the negative result by Abbasi-Yadkori et al. [2013]. On the positive side, it is known from the game theoretic literature that no regret learning would be possible under these conditions if the state space is tree structured as in an extensive form game [Osborne and Rubinstein, 1994]. We leave the study of this improvement, which can be relevant for several games such as Poker, for future works.
Characterizing low concentrability games.
We showed that having access to a queriable expert allows to avoid the concentrability coefficient $\mathcal{C}(\mu^{E},\nu^{E})$. However, expert queries are not always necessary because in some Markov games $\mathcal{C}(\mu^{E},\nu^{E})$ can be upper bounded by a small coefficient. For example, we know from the BC upper bound that the concentrability coefficient equals $1$ when there always exist a unique best response. It is an interesting future direction to study under which setting $\mathcal{C}(\mu^{E},\nu^{E})$ is bounded by a small enough coefficient and therefore we can expect BC to work well.
Appendix D Proofs on BC Upper bound
In this section, we give the omitted proofs for Theorem 3.1. In the first step we state the error decomposition of the Nash Gap
@@ Appendix D Proofs on BC Upper bound
where $\mu^{\star}\in\mathrm{br}(\widehat{\nu})$ and $\nu^{\star}\in\mathrm{br}(\widehat{\mu}).$ We can see that we can split the error into an error for the policy recovered for player 1 depending on the estimation of player 2’s policy ($\mathrm{Error}(\widehat{\nu})$) and for player 1’s policy respectively ($\mathrm{Error}(\widehat{\mu})$). In the following, we will only give the proofs for player 1, as the proofs for player 2 follow analogously. Next, we give a useful lemma that upper-bound the value difference in a two-player game, when one player’s policy is fixed in both value functions, by the total variation. We give the general result and then apply it to our case.
Lemma D.1.
@@
Here, we only prove the first statement. The second statement can be proved by the same idea. By L.1, we have that
[…]
@@
[…]
@@
Applying the result to the BC setting and noting that by definition of the best response we have $V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\mu^{\star},\nu^{\operatorname{E}}}(s_{0})\geq 0$ for $\mu^{\star}\in\mathrm{br}(\widehat{\nu})$ and that the result needs to hold true $\forall\mu\in\mathrm{br}{(\nu)}$, we obtain
@@
Note that the expectation is now over the expert policy, therefore we can use the dataset to get an estimate. Therefore, we apply the standard concentration argument to upper-bound term $\mathbb{E}_{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathrm{TV}\left(\nu^{\operatorname{E}}(\cdot\mid s),\widehat{\nu}(\cdot\mid s)\right)\right]$. By L.2 and union bound, with probability at least $21-\delta/$ 2, $\forall s\in{\mathcal{S}}$
@@
Then we can have that
@@
where in $\mathrm{(i)}$ we applied Cauchy Schwarz and in $\mathrm{(ii)}$ we applied Lemma L.3 and we denoted $N$ as the size of the dataset.
@@
$V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\mu^{\star},\nu^{\operatorname{E}}}(s_{0})\leq\frac{8}{(1-\gamma)^{2}}\max_{\mu\in\mathrm{br}(\widehat{\nu})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{B}}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}$
@@
$V^{\mu^{\operatorname{E}},\nu^{\star}}(s_{0})-V^{\widehat{\mu},\nu^{\star}}(s_{0})\leq\frac{8}{(1-\gamma)^{2}}\max_{\nu\in\mathrm{br}(\widehat{\mu})}\left\|\frac{d^{\mu^{\operatorname{E}},\nu}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}.$
@@
Finally, by using the error decomposition derived in the first step and defining the concentrability coefficient $\mathcal{C}(\hat{\mu},\hat{\nu}):=\max_{\mu\in\mathrm{br}({\widehat{\nu}})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}+\max_{\nu\in\mathrm{br}({\widehat{\mu}})}\left\|\frac{d^{\mu^{\operatorname{E}},\nu}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}$ and use that $\left|{{\mathcal{A}}_{\max}}\right|=\max\{\left|{{\mathcal{A}}}\right|,\left|{{\mathcal{B}}}\right|\}$ we obtain with probability of at least $1-\delta$
@@
$V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\widehat{\mu},\nu^{\star}}(s_{0})\leq{\mathcal{C}}(\widehat{\mu},\widehat{\nu})\frac{8}{(1-\gamma)^{2}}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}$
completing the proof of Theorem 3.1.
Appendix E Proof for necessity of $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$
@@
In this section we give the proof of Theorem 3.2, showing the necessity of a bounded $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ for non-interactive imitation learning even if the learner is fully-aware of the transition model. For a better understanding of the following proof it is essential to remind ourselves of a (simplified) Markov Game hardness construction introduced in Section 3.1 and illustrated in Fig. 1.
Proof of Theorem 3.2.
@@
Let us consider the following family of Zero-Sum Markov Games ${\mathcal{G}}_{\infty}=\{{\mathcal{G}}_{i}\}_{i=1}^{\left|{{\mathcal{A}}}\right|}$ with action spaces of the same cardinality $\left|{{\mathcal{A}}}\right|\geq 3$ given by ${\mathcal{A}}=\{a_{1},a_{2}\ldots,a_{i-1},a_{E},a_{i+1},\ldots,a_{\left|{{\mathcal{A}}}\right|}\}$, =${\mathcal{B}}=\{b_{1},b_{2}\ldots,b_{i-1},b_{E},b_{i+1},\ldots,b_{\left|{{\mathcal{A}}}\right|}\}$ and a shared state space for both agents, given by ${\mathcal{S}}=\{s_{0},s_{1},s_{2},s_{3},s^{\prime}_{3_{1}},\ldots,s^{\prime}_{3_{2\left|{{\mathcal{A}}}\right|-1}},s_{\mathrm{xplt1}_{1}},\ldots,s_{\mathrm{xplt1}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}},s_{\mathrm{xplt2}_{1}},\ldots,s_{\mathrm{xplt2}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}}\}$. From now on we can divide the state space into 5 parts. We have $S_{\mathrm{expert}}:=\{s_{0},s_{2},s_{3}\}$ the states visited by the expert, $s_{1}$ the gating state, $S_{\mathrm{xplt1}}:=\{s_{\mathrm{xplt1}_{1}},\ldots,s_{\mathrm{xplt1}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}}\}$ the states where player 1 can be exploited, $S_{\mathrm{xplt2}}:=\{s_{\mathrm{xplt2}_{1}},\ldots,s_{\mathrm{xplt2}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}}\}$ the states where player 2 can be exploited and $S_{\mathrm{copy}}:=\{s^{\prime}_{3_{1}},\ldots,s^{\prime}_{3_{2\left|{{\mathcal{A}}}\right|-1}}\}$ that are copies of the final states visited by the expert in the sense that they are sharing the same reward. Additionally, consider a dirac on state $s_{0}$ as an initial state distribution, i.e. $d_{0}(s)=\delta_{s_{0}}$. The transition model is deterministic and in all states except $s_{0}$ and $s_{1}$, simply a transition to one neighboring state. Therefore, we only give a detailed description for these two states. And for the gating state $s_{1}$, we only consider the next potential set of states, as states inside these sets share the same reward function and each action pair leads to a unique state inside these sets, i.e. every action combination leads to a different state. In particular, we have
@@
[…]
@@
An illustration of the described Markov Game for $\left|{{\mathcal{A}}}\right|=\left|{{\mathcal{B}}}\right|=3$ can be found in Fig. 1, where $a_{i}=a_{3}$ and $b_{i}=b_{3}$.
@@
Now we show that for the family of Markov Games ${\mathcal{G}}_{\infty}$ where for each ${\mathcal{G}}_{i}\in{\mathcal{G}}_{\infty}$ it holds that ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$, any learning algorithm $\mathrm{Alg}$ has a Nash Gap of the order $(1-\gamma)^{-1}$. For that let $(\widehat{\mu},\widehat{\nu})$ be the output of any non-interactive imitation learning algorithm $\mathrm{Alg}$ with data from $(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$. It is important to observe that since all games in ${\mathcal{G}}_{\infty}$ are identical in $s_{0}$ and they differs only in transition dynamics from $s_{1}$. However, no information about $s_{1}$ is available in $\mathcal{D}$. Therefore, the learner has no mean to distinguish which game she is facing. For this reason $\widehat{\mu},\widehat{\nu}$ do not depend on the game index $i$. Then denoting by $A_{\mathrm{Alg}}$ and $B_{\mathrm{Alg}}$ the action played by the learner in the state $s_{1}$, it holds true that
@@
=$\overset{(i)}{=}\frac{1}{(1-\gamma)}\left(\frac{\sum_{i=1}^{\left|{{\mathcal{A}}}\right|}\mathbb{P}(B_{\mathrm{Alg}}\neq b_{i})}{\left|{{\mathcal{A}}}\right|}+\frac{\sum_{i=1}^{\left|{{\mathcal{A}}}\right|}\mathbb{P}(A_{\mathrm{Alg}}\neq a_{i})}{\left|{{\mathcal{A}}}\right|}\right)$
@@
Additionally, note that even if the learner has access to the transition dynamics, the learner can not differentiate the actions from $s_{1}$ as all actions lead to different states. Therefore, she cannot use this knowledge to recover an action that would lead to $s\in S_{\mathrm{copy}}$, which would avoid a regret of the order$(1-\gamma)^{-1}$. This completes the proof.
Appendix F Analysis of BR Oracle Algorithm
This section presents the analysis of Algorithm 1 which provides a sample complexity guarantee without requiring single deviation concentrability under the assumption of a Best Response Oracle (Definition 4.1). The difference to the later presented algorithm MURMAIL Algorithm 2 is that here we can query the best response oracle to sample from the expectation that contains the best response of the current policy $\mu_{k}$ and $\nu_{k}$ respectively. In particular, we sample a state from the induced discounted state distributions. However, as noted in our discussion around Eq. 3 we first have to transform the optimization problem into a convex one as the original objective of our optimization problem is non-convex. This results in Eq. 4, from where we can obtain a Martingale difference sequence and a regret term, which then can be minimized as we now can construct an unbiased gradient estimator and use a version of online mirror descent to construct an update of our policies.

### v2

@@ 6 Conclusion and Future Directions
This paper provides the first sample complexity analysis of behavioral cloning in the multi-agent setting. The provided upper bound depends on the all policy deviation concentrability coefficient. Additionally, which is shown to be unavoidable in general. Unfortunately, it is quite easy to come up with MGs where the concentrability coefficient is unbounded. In particular, we introduce the smaller quantity ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is unbounded and where no non-interactive MAIL algorithm can learn an equilibrium from data. In this situation, we resort to expert queries and we introduce novel algorithms dubbed MAIL-BRO and MURMAIL, which achieve an $\varepsilon$-approximate Nash equilibrium with a polynomial number of expert queries and computational cost polynomial in all problem parameters.
We outline a few interesting future directions and research questions left open by our work.
Lower Bound.
This work has provided a hardness result, showing that if ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is unbounded, there exists no non-interactive MAIL algorithm that can learn. However, it holds true that ${\mathcal{C}}_{\max}\geq{\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ and it remains an open question if there exists a Markov Game, where ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is bounded but ${\mathcal{C}}_{\max}=\infty$ and no non-interactive MAIL algorithm can recover an equilibrium. Answering this would close the gap between the provided upper bound of BC and the lower bound. In particular, this would answer what the fundamental quantity of non-interactive MAIL is. Additionally, it remains an open question if one can prove a statistical lower bound featuring ${\mathcal{C}}({\mu^{\operatorname{E}},\nu^{\operatorname{E}}})$ or ${\mathcal{C}}_{\max}$ and whether BC is rate optimal in $\varepsilon.$
The focus of this work was to show the first sample complexity bound for a computationally efficient algorithm in the queriable expert setting. For the sake of simplicity, we did not try to optimize the dependence of the upper bound in the accuracy parameters $\varepsilon$, effective horizon $(1-\gamma)^{-1}$, states and actions cardinality $\left|{{\mathcal{S}}}\right|$ and $\left|{{\mathcal{A}}}\right|$. A possible direction of improvement is to derive a tighter analysis of the outer loop using faster rates for the regret of the squared loss (see for example (Cesa-Bianchi and Lugosi, 2006, Chapter 3)).
Moreover, the upper bound could be improved by removing the need of the RL inner loop in MURMAIL. In general, replacing it with a no regret learner that minimizes the regret in an MDP with changing reward function and transitions is not possible because of the negative result by Abbasi-Yadkori et al. (2013). On the positive side, it is known from the game theoretic literature that no regret learning would be possible under these conditions if the state space is tree structured as in an extensive form game (Osborne and Rubinstein, 1994). We leave the study of this improvement, which can be relevant for several games such as Poker, for future works.
Extension to deep imitation learning.
The current analysis is limited to tabular Markov Games. However, the main conceptual ideas easily carry on to deep imitation learning experiments. The largest theory-practice gap would be in the inner loop where UCBVI would need to be replaced by a Deep RL algorithm such as DQN Mnih et al. (2015) or Soft Actor Critic Haarnoja et al. (2018), just to name a few.
@@ Appendix A Notation
$\varepsilon$-NE $\varepsilon$-approximate Nash Equilibrium
@@
NN Number of trajectories in dataset
@@
[…]
@@
$\varepsilon_{opt}$ Optimality gap for RL inner loop
@@ Appendix B Related Work
Strikingly, the scenario is completely swapped in the multi-agent setting. Our negative result Theorem˜3.2 shows that given knowledge of the transition no significant improvements over BC can be expected. On the other side, our positive result Theorem˜4.2 shows that in the interactive setting a consistent improvement is expected over BC if $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is large.
@@
Appendix C Proofs on BC Upper bound
In this section, we give the omitted proofs for Theorem˜3.1. In the first step we state the error decomposition of the Nash Gap
@@ Appendix C Proofs on BC Upper bound
[…]
Lemma C.1.
@@
Here, we only prove the first statement. The second statement can be proved by the same idea. By K.1, we have that
$\begin{aligned} V^{\mu,\nu}(s_{0})-V^{\mu,\nu^{\prime}}(s_{0})&=\mathbb{E}_{\mu,\nu}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\mathbb{E}_{(a,b)\sim(\mu,\nu)}\left[Q^{\mu,\nu^{\prime}}(s,a,b)\right]-\mathbb{E}_{(a,b)\sim(\mu,\nu^{\prime})}\left[Q^{\mu,\nu^{\prime}}(s,a,b)\right]\right)\right].\end{aligned}$
@@
$\begin{aligned} \left|V^{\mu,\nu}(s_{0})-V^{\mu,\nu^{\prime}}(s_{0})\right|\leq\mathbb{E}_{\mu,\nu}\left[\sum_{t=0}^{\infty}\gamma^{t}\left|\mathbb{E}_{(a,b)\sim(\mu,\nu)}\left[Q^{\mu,\nu^{\prime}}(s,a,b)\right]-\mathbb{E}_{(a,b)\sim(\mu,\nu^{\prime})}\left[Q^{\mu,\nu^{\prime}}(s,a,b)\right]\right|\right].\end{aligned}$
@@
[…]
@@
Note that we can do the change of measure also with respect to a more general offline distribution $\rho.$ Here, we use $\rho=d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}.$ If we would use the general offline sampling distribution, we would have
$V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\mu^{\star},\nu^{\operatorname{E}}}(s_{0})$ $=\frac{2}{1-\gamma}\max_{\mu\in\mathrm{br}(\widehat{\nu})}\sum_{t=0}^{\infty}\gamma^{t}\sum_{s\in{\mathcal{S}}}\frac{d^{\mu,\nu^{\operatorname{E}}}(s)}{\rho(s)}\rho(s)\mathrm{TV}\left(\nu^{\operatorname{E}}(\cdot\mid s),\widehat{\nu}(\cdot\mid s)\right).$
$\leq\frac{2}{1-\gamma}\max_{\mu\in\mathrm{br}({\widehat{\nu}})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{\rho}\right\|_{\infty}\mathbb{E}_{\rho}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathrm{TV}\left(\nu^{\operatorname{E}}(\cdot\mid s),\widehat{\nu}(\cdot\mid s)\right)\right].$
[…]
$\frac{2}{1-\gamma}\max_{\mu\in\mathrm{br}({\widehat{\nu}})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\mathbb{E}_{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathrm{TV}\left(\nu^{\operatorname{E}}(\cdot\mid s),\widehat{\nu}(\cdot\mid s)\right)\right]$
$\leq\frac{2}{1-\gamma}\max_{\nu\in\Pi}\max_{\mu\in\mathrm{br}({\nu})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\mathbb{E}_{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathrm{TV}\left(\nu^{\operatorname{E}}(\cdot\mid s),\widehat{\nu}(\cdot\mid s)\right)\right],$
where $\nu\in\Pi$ is any admissible policy for the second player. Taking the maximum over the best responses to this more general policy class forms an upper bound.
Additionally, note that the expectation is now over the expert policy, therefore we can use the dataset to get an estimate. Therefore, we apply the standard concentration argument to upper-bound term $\mathbb{E}_{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathrm{TV}\left(\nu^{\operatorname{E}}(\cdot\mid s),\widehat{\nu}(\cdot\mid s)\right)\right]$. By K.2 and union bound, with probability at least $1-\delta/2$, $\forall s\in{\mathcal{S}}$
@@
Then, we can have that
@@
where in $\mathrm{(i)}$ we applied Cauchy Schwarz and in $\mathrm{(ii)}$ we applied Lemma˜K.3 and we denoted $N$ as the size of the dataset.
@@
$V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\mu^{\star},\nu^{\operatorname{E}}}(s_{0})\leq\frac{8}{(1-\gamma)^{2}}\max_{\nu\in\Pi}\max_{\mu\in\mathrm{br}(\nu)}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{B}}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}$
@@
$V^{\mu^{\operatorname{E}},\nu^{\star}}(s_{0})-V^{\widehat{\mu},\nu^{\star}}(s_{0})\leq\frac{8}{(1-\gamma)^{2}}\max_{\mu\in\Pi}\max_{\nu\in\mathrm{br}(\mu)}\left\|\frac{d^{\mu^{\operatorname{E}},\nu}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}.$
@@
Finally, by using the error decomposition derived in the first step, defining $\mathcal{C}_{\max}:=\max\left\{\max_{\nu\in\Pi}\max_{\mu\in\mathrm{br}({\nu})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{\rho}\right\|_{\infty},\,\max_{\mu\in\Pi}\max_{\nu\in\mathrm{br}({\mu})}\left\|\frac{d^{\mu^{\operatorname{E}},\nu}}{\rho}\right\|_{\infty}\right\}$ with $\rho=d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}$ and use that $\left|{{\mathcal{A}}_{\max}}\right|=\max\{\left|{{\mathcal{A}}}\right|,\left|{{\mathcal{B}}}\right|\}$ we obtain with probability of at least $1-\delta$
@@
$V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\widehat{\mu},\nu^{\star}}(s_{0})\leq{\mathcal{C}}_{\max}\frac{8}{(1-\gamma)^{2}}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}$
completing the proof of Theorem˜3.1.
Appendix D Proof for necessity of $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$
@@
In this section we give the proof of Theorem˜3.2, showing the necessity of a bounded $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ for non-interactive imitation learning even if the learner is fully-aware of the transition model. For a better understanding of the following proof it is essential to remind ourselves of a (simplified) Markov Game hardness construction introduced in Section˜3.1 and illustrated in Fig.˜1.
Proof of Theorem˜3.2.
@@
Let us consider the following family of Zero-Sum Markov Games ${\mathcal{G}}_{\infty}=\{{\mathcal{G}}_{i}\}_{i=1}^{\left|{{\mathcal{A}}}\right|}$ with action spaces of the same cardinality $\left|{{\mathcal{A}}}\right|\geq 3$ given by ${\mathcal{A}}=\{a_{1},a_{2}\ldots,a_{i-1},a_{E},a_{i+1},\ldots,a_{\left|{{\mathcal{A}}}\right|}\}$, ${\mathcal{B}}=\{b_{1},b_{2}\ldots,b_{i-1},b_{E},b_{i+1},\ldots,b_{\left|{{\mathcal{A}}}\right|}\}$ and a shared state space for both agents ${\mathcal{S}}=\{s_{0},s_{1},s_{2},s_{3},s^{\prime}_{3_{1}},\ldots,s^{\prime}_{3_{2\left|{{\mathcal{A}}}\right|-1}},s_{\mathrm{xplt1}_{1}},\ldots,s_{\mathrm{xplt1}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}},s_{\mathrm{xplt2}_{1}},\ldots,s_{\mathrm{xplt2}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}}\}$. From now on we can divide the state space into 5 parts. We have $S_{\mathrm{expert}}:=\{s_{0},s_{2},s_{3}\}$ the states visited by the expert, $s_{1}$ the gating state, $S_{\mathrm{xplt1}}:=\{s_{\mathrm{xplt1}_{1}},\ldots,s_{\mathrm{xplt1}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}}\}$ the states where player 1 can be exploited, $S_{\mathrm{xplt2}}:=\{s_{\mathrm{xplt2}_{1}},\ldots,s_{\mathrm{xplt2}_{(\left|{{\mathcal{A}}}\right|-1)^{2}/2}}\}$ the states where player 2 can be exploited and $S_{\mathrm{copy}}:=\{s^{\prime}_{3_{1}},\ldots,s^{\prime}_{3_{2\left|{{\mathcal{A}}}\right|-1}}\}$ that are copies of the final states visited by the expert in the sense that they are sharing the same reward. Additionally, consider a dirac on state $s_{0}$ as an initial state distribution, i.e. $d_{0}(s)=\delta_{s_{0}}$. The transition model is deterministic and in all states except $s_{0}$ and $s_{1}$, simply a transition to one neighboring state. Therefore, we only give a detailed description for these two states. And for the gating state $s_{1}$, we only consider the next potential set of states, as states inside these sets share the same reward function and each action pair leads to a unique state inside these sets, i.e. every action combination leads to a different state. In particular, we have
@@
$P_{{\mathcal{G}}_{i}}(\cdot\mid s_{1},a,b)\;=\;\begin{dcases*}S_{\mathrm{copy}},&if $(a,b)\in\{(a_{i},b_{i})\}\,\cup\,\{(a_{i},b_{j}),(a_{j},b_{i})\,\mid\,\forall j\neq i\}$,\\[6.0pt] S_{\mathrm{xplt1}},&if $(a,b)\in\mathcal{E}_{p,q}$\\ S_{\mathrm{xplt2}},&otherwise,\end{dcases*}$
@@
An illustration of the described Markov Game for $\left|{{\mathcal{A}}}\right|=\left|{{\mathcal{B}}}\right|=3$ can be found in Fig.˜1, where $a_{i}=a_{3}$ and $b_{i}=b_{3}$.
@@
[…]
@@
$\overset{(i)}{=}\frac{1}{(1-\gamma)}\left(\frac{\sum_{i=1}^{\left|{{\mathcal{A}}}\right|}\mathbb{P}(B_{\mathrm{Alg}}\neq b_{i})}{\left|{{\mathcal{A}}}\right|}+\frac{\sum_{i=1}^{\left|{{\mathcal{A}}}\right|}\mathbb{P}(A_{\mathrm{Alg}}\neq a_{i})}{\left|{{\mathcal{A}}}\right|}\right)$
@@
[…]
Appendix E Analysis of BR Oracle Algorithm
This section presents the analysis of Algorithm˜1 which provides a sample complexity guarantee without requiring single deviation concentrability under the assumption of a Best Response Oracle (Definition˜4.1). The difference to the later presented algorithm MURMAIL Algorithm˜2 is that here we can query the best response oracle to sample from the expectation that contains the best response of the current policy $\mu_{k}$ and $\nu_{k}$ respectively. In particular, we sample a state from the induced discounted state distributions. However, as noted in our discussion around Eq.˜3 we first have to transform the optimization problem into a convex one as the original objective of our optimization problem is non-convex. This results in Eq.˜4, from where we can obtain a Martingale difference sequence and a regret term, which then can be minimized as we now can construct an unbiased gradient estimator and use a version of online mirror descent to construct an update of our policies.